How do you solve? -31=8-6p-7p
Found 2 solutions by stanbon, checkley79:
Answer by stanbon(75887)

-31=8-6p-7p
----
Simplify:
-31 = 8-13p
---
Subtract 8 from both sides:
-39 = -13p
----
Divide both sides by -13 to get:
p = 3
===========
Cheers,
Stan H.
===========

Answer by checkley79(3341)

-31=8-6p-7p
-31-8=-13P
-39=-13P
P=-39/-13
P=3 ANS.